Plz answer: If 1<=a<=2, then will the following be true? sqrt [a+2sqrt(a-1)] + sqrt [a-2sqrt(a-1)] = 2 Thanks.

Plz answer:
 
If 1<=a<=2, then will the following be true?
 
sqrt [a+2sqrt(a-1)] + sqrt [a-2sqrt(a-1)] = 2
 
Thanks.

Grade:12th pass

1 Answers

Nidhi Singh
5 years ago
Yes it holds.Sqrt (a+2sqrt(a-1))+sqrt(a-2sqrt(a-1))=2Squaring both sides,a+2sqrt(a-1)+a-2sqrt ( a-1)+ 2sqrt[(a+2sqrt(a-1))(a-2sqrt (a-1))]=42a+2sqrt{a^2-4a+4}=4a(+-)(a-2)=2Taking +ve value satisfies as a=2Taking-ve value satisfies as 2=2